Jacket$65
Discount$30%
Sales Tax$4%

Answer by queenofit(120)

you have a jacket that cost 65 and get a discount of 30% plus the sales tax of 4%. Ok, first let me tell you that to get the cost of the jacket most people will multiply 65 * .30 then subtract the answer from the 65 to get the cost of the jacket after the discount. I will show that on the left, however, I like to save a step by multiplying by 70%. there is a 100% in everything, 100-30 gives you 70%. You are payiing 70% of the origional cost, so I multiply by this number and I don't have to subtract the answer from the cost of the jacket. I will show this on the right. When you get your cost after the discount, you have to add the sales tax of 4%. Remember to move your decimal two places to the left when changing from a percent to a decimal.
65*.30=19.50 (discount) 65*.70=45.50 (cost after discount)
65-19.50=45.50 (your cost after discount) 45.50*.04=1.82
45.50*.04=1.82 (tax) 45.50+1.82=47.32
45.50+1.82=47.32(your final cost)